A Bachelor of Commerce student has to take 8 courses to graduate. If none of the courses are prerequisites for others, how many groups of 3 courses can the student select for the coming semester?
1
Expert's answer
2020-11-16T07:34:00-0500
"C_8^3=\frac {8!}{3!(8-3)!}=56"